# New properties of the $\varphi$-representation of integers

### v1

@@ Abstract
We prove a few new properties of the $\varphi$-representation of integers, where $\varphi=(1+\sqrt{5})/2$. In particular, we prove a 2012 conjecture of Kimberling.
@@ 1 Introduction
Let $\varphi=(1+\sqrt{5})/2$ be the golden ratio. Bergman [1] showed that every non-negative real number $x$ has a representation as a sum of distinct (possibly negative) powers of $\varphi$; we write $x=\sum_{i\leq t}e_{i}\varphi^{i}$ for $e_{i}\in\{0,1\}$. Furthermore, this representation is unique, provided that the representation obeys two rules:
  1. 1.
    $e_{i}e_{i+1}=0$ for all $i$;
  2. 2.
    there is no integer $j$ such that $e_{2i}=1$, $e_{2i-1}=0$ for all $i\leq j$.
@@
If the $\varphi$-representation is finite, we can write it as a binary string with a decimal point, in the form $e_{t}\cdots e_{1}e_{0}.e_{-1}e_{-2}e_{-3}\cdots$ representing $\sum_{i\leq t}e_{i}\varphi^{i}$. Thus, for example, $5$ is represented by the binary string $1000.1001$.
It is known [17] that the $\varphi$-representation of every natural number consists of only a finite number of powers of $\varphi$. In fact, this property of a non-negative real number xx—that is, having a finite $\varphi$-representation—is exactly the same as being an element of $\mathbb{Z}[\varphi]$. In this note we establish some novel properties of the $\varphi$-representations of natural numbers. We use both “conventional” arguments and the Walnut theorem-prover.
@@ 2 Shevelev’s set and Kimberling’s conjecture
First, suppose $x$ is an integer whose $\varphi$-expansion consists of all even exponents. Let ${\cal E}_{\text{pos}}$ be the set of positive exponents and ${\cal E}_{\text{neg}}$ be the set of negative exponents. Consider the number $n$ whose $\varphi$-representation consists of powers of $\varphi$ in ${\cal E}_{\text{pos}}$ and ${\cal E}_{\text{pos}}$. By Lemma 2 we see that $n$ is an integer. Furthermore $0\leq|n-x|<\varphi^{-2}+\varphi^{-4}+\varphi^{-6}+\cdots=\varphi^{-1}<1$. Since $n$ is an integer, we must have $n=x$ and its expansion consists of the exponents ${\cal E}_{\text{pos}}$, ${\cal E}_{\text{pos}}$, and possibly 0. So it is antipalindromic.
@@
The set $S$ consists of those natural numbers $n$ such that $\sum_{i\leq t}e_{i}(n)\varphi^{2i}$ is an integer, where $n=\sum_{i\leq t}e_{i}(n)\varphi^{i}$ is the $\varphi$-representation of $n$.
@@
Let $\sum_{i\leq t}e_{i}\varphi^{i}$ be the $\varphi$-representation of an integer $n$, and define $x=\sum_{i\leq t}e_{i}\varphi^{2i}$.
@@ 3 Other results
To explain what Walnut can do and how it does it, we need to talk about the Zeckendorf (or Fibonacci) representation of natural numbers. In this system [12, 21] we write a natural number $n$ as a sum of distinct Fibonacci numbers $F_{i}$ for $i\geq 2$. This representation is unique, provided we never use two adjacent Fibonacci numbers in it. We abbreviate such a representation as a binary string $b_{t}b_{t-1}\cdots b_{2}$ for the integer $b_{t}b_{t-1}\cdots b_{2}]_{F}:=\sum_{2\leq i\leq t}b_{i}F_{i}$. Thus, for example, $11=F_{6}+F_{4}$ is represented by the binary string $10100$.
In the paper [20], the author showed how to construct a finite automaton $\tt saka$} that takes three binary strings in parallel as input, say $w,y,z$, and accepts if $w]_{F}$ has $\varphi$ representation $y.z^{R}$, where $z^{R}$ denotes the reversal of the string $z$. Although the automaton saka takes binary strings as input, in Walnut we can call the automaton with numerical inputs, and sometimes this is quite useful. In that case, we use the Zeckendorf representation of integers. For example, if we write
eval check "?msd_fib \$saka(5,5,6)":
in Walnut, this is asking whether $1000.1001$ is the $\varphi$-representation of $5$, and Walnut answers TRUE.
@@
We have seen that the set $S$ consists of those integers whose $\varphi$-representation consists of all even powers. This raises the question of whether there are any integers whose $\varphi$-representation consists of all odd powers. This is answered (negatively) in the following result.
@@
The smallest exponent of the $\varphi$-representation of a positive integer must be even. Furthermore, if $n\geq 2$ and $L_{2i-1}<n\leq L_{2i+1}$, then the smallest exponent is 2i-2i.
@@
We use Walnut. For the first claim, we use
@@
eval small_test "?msd_fib An,x,y (n>=2 & \$saka(n,x,y)) => \$largesteven(y)":
@@
For the second claim, we recall the formula $L_{n}=F_{n-1}+F_{n+1}$ for $n\geq 1$, as well as the equalities $L_{0}=F_{3}$ and $L_{2}=F_{4}$. We use the following code:
@@
The last expression returns TRUE; it checks that if $L_{2i-1}<n\leq L_{2i+1}$, then $\varphi^{-2i}$ is the largest power of $\varphi$ appearing in the $\varphi$-representation $x.y^{R}$ of $n$. This means that the leftmost $1$ in the string $y^{R}$ corresponds to $F_{2i+1}$, if we considered $y^{R}$ as a Zeckendorf expansion. So the largest $1$ digit in $y^{R}$, when viewed as a Zeckendorf representation of $n$, corresponds to $F_{2i+1}$, which is the meaning of the implicand. ∎
We also provide a more standard proof. The proof of the first claim was shown to us by ChatGPT 5.
@@
Suppose $n\geq 1$ is an integer where the least exponent $m$ appearing in the $\varphi$-representation is odd. Then $m+1$ does not appear. Apply Galois conjugation, replacing $\varphi$ in the representation by $\overline{\varphi}$, obtaining that $n$ is also a sum of the same powers of $\overline{\varphi}$. Now since $m$ is odd, we have $\overline{\varphi}^{m}<0$. Thus the next larger positive term one could have in this new representation is $\overline{\varphi}^{m+3}$. It follows that $n\leq\overline{\varphi}^{m}+\overline{\varphi}^{m+3}+\overline{\varphi}^{m+5}+\cdots=\overline{\varphi}^{m}(1-\varphi^{-2})<0$, a contradiction.
For the second claim let $n\geq 2$ be an integer. By the first part we know that the smallest exponent in its $\varphi$-representation is even, say 2i-2i. We need to show that $L_{2i-1}<n\leq L_{2i+1}$. By Galois conjugation we have $n=\varphi^{2i}\pm\dots$, where the remaining terms have exponents $\leq 2i-2$ and all even powers occur with positive sign and all odd powers occur with negative sign. Thus we get on the one hand
@@
  • $n$ has one non-negative even exponent and no negative even exponents. If we say that the last position of $x$ is “position 11”, then this corresponds to $x$ having exactly one $1$ at an odd position (and even positions don’t matter) and $y$ having no $1$’s at even positions (and odd positions don’t matter).
@@
def thm7 "?msd_fib Ex,y \$saka(n,x,y) & \$one_even_exponent(x,y)":
@@
$\varphi^{2k}-\varphi^{-(2i+1)}=\varphi^{-(2k+1)}+\dots+\varphi^{-(2i-1)}+\varphi^{-(2i+2)}.$
@@
Now assume conversely that $x>1$ has exactly one even exponent in its $\varphi$-representation. By Theorem 6 it must be the smallest exponent, and so in particular all positive exponents are odd, and $\varphi^{0}$ does not appear. Let ${\cal O}_{\text{pos}}$ be the set of (odd) positive exponents in the $\varphi$-representation of $x$, and ${\cal A}_{\text{neg}}$ be the set of negative exponents. Consider the number
@@
eval thm8b "?msd_fib An \$sum_even_lucas(n-2) <=> \$thm8(n)":
@@
eval thm8c "?msd_fib An \$thm8(n) => Ex,y \$saka(n,x,y) & \$fibmatch(x,2)":
@@
def thm9_match "?msd_fib \$fib3(a,d) & \$fibmatch(x,d) & \$fibmatch(y,a)":
eval thm9b "?msd_fib An \$thm9(n) =>  Ex,y \$saka(n,x,y) & ((\$fibmatch(x,5)
   & \$fibmatch(x,2)) | Ea,d \$thm9_match(a,d,x,y))":

### v2

@@ Abstract
We prove a few new properties of the $\varphi$-representation of integers, where $\varphi=(1+\sqrt{5})/2$. In particular, we prove a 2012 conjecture of Kimberling. As software assistants, we used the Walnut theorem-prover, and in one proof, ChatGPT 5.
@@ 1 Introduction
Let $\varphi=(1+\sqrt{5})/2$ be the golden ratio. Bergman [1] showed that every positive integer $x$ can be represented as a finite sum of distinct (possibly negative) powers of $\varphi$; we write $x=\sum_{a\leq i\leq b}e_{i}\varphi^{i}$ for $a,b\in\mathbb{Z}$ with $a\leq b$ and $e_{i}\in\{0,1\}$. Furthermore, this representation is unique, provided that $e_{i}e_{i-1}=0$ for $a+1\leq i\leq b$. Also see [17].
More generally, Rényi [16] proved an analogous result for representing arbitrary positive real numbers. Here the representation could be infinite, and to get uniqueness we have to impose the further rule that there is no index $j$ such that $e_{j-2i}=1$, $e_{j-2i-1}=0$ for all $i\geq 0$.
@@
If a $\varphi$-representation is finite, we can write it as a binary string with a decimal point, in the form $e_{b}\cdots e_{1}e_{0}.e_{-1}e_{-2}\cdots e_{a}$ representing $\sum_{a\leq i\leq b}e_{i}\varphi^{i}$. If $e_{b}\not=0$ and $e_{a}\not=0$, we call this representation canonical. We call $e_{b}\cdots e_{1}e_{0}$ the positive part and $e_{-1}e_{-2}\cdots e_{a}$ the negative part, referring to the exponents that appear. Thus, for example, the canonical representation of $5$ is the string $1000.1001$.
Finite $\varphi$-representations correspond exactly to the non-negative members of $\mathbb{Z}[\varphi]$. This was mentioned in Bergman [1] and generalized in [11]. In this note we establish some novel properties of the $\varphi$-representations of natural numbers. We use both “conventional” arguments and the Walnut theorem-prover.
@@ 2 Shevelev’s set and Kimberling’s conjecture
First, suppose $x$ is an integer whose $\varphi$-expansion consists of all even exponents. Let ${\cal E}_{\text{pos}}$ be the set of positive exponents and ${\cal E}_{\text{neg}}$ be the set of negative exponents. Consider the number $n$ whose $\varphi$-representation consists of powers of $\varphi$ in ${\cal E}_{\text{pos}}$ and ${\cal E}_{\text{pos}}$, and possibly $\{0\}$ if 0 is present as an exponent in the expansion of $x$. By Lemma 2 we see that $n$ is an integer. Furthermore $0\leq|n-x|<\varphi^{-2}+\varphi^{-4}+\varphi^{-6}+\cdots=\varphi^{-1}<1$. Since $n$ is an integer, we must have $n=x$ and its expansion consists of the exponents ${\cal E}_{\text{pos}}$, ${\cal E}_{\text{pos}}$, and possibly 0. So it is antipalindromic.
@@
The set $S$ consists of those natural numbers $n$ such that $\sum_{a\leq i\leq b}e_{i}(n)\varphi^{2i}$ is an integer, where $n=\sum_{a\leq i\leq b}e_{i}(n)\varphi^{i}$ is the $\varphi$-representation of $n$.
@@
Let $\sum_{a\leq i\leq b}e_{i}\varphi^{i}$ be the $\varphi$-representation of an integer $n$, and define $x=\sum_{a\leq i\leq b}e_{i}\varphi^{2i}$.
@@ 3 Other results
To explain what Walnut can do and how it does it, we need to talk about the Zeckendorf (or Fibonacci) representation of natural numbers. In this system [12, 21] we write a natural number $n$ as a sum of distinct Fibonacci numbers $F_{i}$ for $i\geq 2$. This representation is unique, provided we never use two adjacent Fibonacci numbers in it. We abbreviate such a representation as a binary string $b_{t}b_{t-1}\cdots b_{2}$ for the integer $b_{t}b_{t-1}\cdots b_{2}]_{F}:=\sum_{2\leq i\leq t}b_{i}F_{i}$; it is called canonical if $b_{t}=1$. Thus, for example, $11=F_{6}+F_{4}$ is represented by the binary string $10100$.
In the paper [20], the first author showed how to construct a $39$-state finite automaton $\tt saka$} that takes three same-length binary strings in parallel as input, say $w,y,z$, and accepts if $w]_{F}$ has $\varphi$ representation $y.z^{R}$, where $z^{R}$ denotes the reversal of the string $z$. (The automaton saka can be downloaded from https://cs.uwaterloo.ca/~shallit/Papers/allcode.tar.) In order to force the strings to have the same length, we pad the shorter of $w$, $y$, and $z$ with leading zeroes, if needed. Furthermore, any number of leading zeroes is allowed. Another way to think of this representation $y.z^{R}$ is as a concatenation of the blocks $[0,0],[0,1],[1,0],[1,1]$, where the first components spell out $y$ and the second components spell out $z$. We call this the folded representation.
Although the automaton saka takes binary strings as input, in Walnut we can call the automaton with numerical inputs, and sometimes this is quite useful. For example, if we write
eval check "?msd_fib \$saka(9,10,7)":
in Walnut, this is asking whether $10010.0101$ is the $\varphi$-representation of $9$ (since $10]_{F}=10010$ and $7]_{F}=1010$). Here Walnut answers TRUE.
@@
Example 6.
The following Walnut code generates an automaton accepting those pairs $(x,y)$ such that $x.y^{R}$ is the $\varphi$-representation of some integer $n\geq 0$. The automaton is depicted in Figure 1.
def sk "?msd_fib En \$saka(n,x,y)":
Figure 1: Automaton accepting $(x,y)$ such that $x.y^{R}$ is the $\varphi$-representation of some integer $n\geq 0$.
As a warmup let us prove the following result.
Proposition 7.
Let $n$ be a positive integer with canonical representation $\sum_{a\leq i\leq b}e_{i}\varphi^{i}$. If $b$ is odd then a=(b+1)a=-(b+1), and if $b$ is even then a=ba=-b.
Proof.
In terms of the folded representation this means that for $n\geq 1$, an even-length representation must begin with $[1,1]$, and for $n\geq 2$ an odd-length representation must begin with $[1,0][0,1]$. We can verify this with Walnut as follows:
reg start11 msd_fib msd_fib "[0,0]*[1,1]([0,0]|[0,1]|[1,0]|[1,1])*":
reg start1001 msd_fib msd_fib "[0,0]*[1,0][0,1]([0,0]|[0,1]|[1,0]|[1,1])*":
reg evenl msd_fib msd_fib "[0,0]*([0,1]|[1,0]|[1,1])
(([0,0]|[0,1]|[1,0]|[1,1])([0,0]|[0,1]|[1,0]|[1,1]))*([0,0]|[0,1]|[1,0]|[1,1])":
reg oddl msd_fib msd_fib "[0,0]*([0,1]|[1,0]|[1,1])
(([0,0]|[0,1]|[1,0]|[1,1])([0,0]|[0,1]|[1,0]|[1,1]))*":
eval check_num_even "?msd_fib An,x,y (n>=1 & \$saka(n,x,y) & \$evenl(x,y)) =>
\$start11(x,y)":
eval check_num_odd "?msd_fib An,x,y (n>=2 & \$saka(n,x,y) & \$oddl(x,y)) =>
\$start1001(x,y)":
And Walnut returns TRUE for both. ∎
We have seen that the set $S$ consists of those integers $n$ whose $\varphi$-representation consists of all even powers. We can find an automaton accepting the set $S$ as follows:
reg shiftr {0,1} {0,1} "([0,0]|[1,0][1,1]*[0,1])*(()|[1,0][1,1]*)":
# accepts x,y such that y is the right shift of the bits of x
def a178482 "?msd_fib Ex,y \$saka(n,x,y) & \$shiftr(x,y)":
It is depicted in Figure 2.
Figure 2: Automaton accepting the Zeckendorf representations of the members of sequence A178482.
This raises the question of whether there are any integers whose $\varphi$-representation consists of all odd powers. This is answered (negatively) in the following version of a theorem of Dekking [6, Thm. 9]. It says that the negative parts of canonical $\varphi$-representations consist of all binary strings without two consecutive $1$’s that end with a $1$ and whose length is even.
@@
  • (a)
    Suppose $n\geq 2$ is a positive integer with canonical $\varphi$-representation $x.y$ for strings $x,y$. Then $|y|$ is positive and even.
  • (b)
    Let $y$ be the canonical Zeckendorf representation of some positive integer such that $|y|$ is even. Then there exists an integer $n$ and a string $x$ such that the $\varphi$-representation of $n$ is $n=x.y^{R}$.
  • (c)
    If $n\geq 2$ and $L_{2i-1}<n\leq L_{2i+1}$, then the smallest exponent in the $\varphi$-representation of $n$ is 2i-2i.
@@
We use Walnut. For claim (a), we use the following code:
@@
eval thm_a "?msd_fib An,x,y (n>=2 & \$saka(n,x,y)) => \$largesteven(y)":
@@
For claim (b), we write
eval thm_b "?msd_fib Ay Ez (\$largesteven(y) & y=z) => En,x \$saka(n,x,z)":
and once again we get TRUE.
For claim (c), we recall the formula $L_{n}=F_{n-1}+F_{n+1}$ for $n\geq 1$, as well as the equalities $L_{0}=F_{3}$ and $L_{2}=F_{4}$. We use the following code:
@@
The last expression returns TRUE; it checks that if $L_{2i-1}<n\leq L_{2i+1}$, then $\varphi^{-2i}$ is the least power of $\varphi$ appearing in the $\varphi$-representation $x.y^{R}$ of $n$. This means that the leftmost $1$ in the string $y^{R}$ corresponds to $F_{2i+1}$, if we considered $y^{R}$ as a Zeckendorf expansion. So the largest $1$ digit in $y^{R}$, when viewed as a Zeckendorf representation of $n$, corresponds to $F_{2i+1}$, which is the meaning of the implicand. ∎
We also provide a more standard, non-Walnut proof. The proof of the first claim was shown to us by ChatGPT 5.
@@
(a) Suppose $n\geq 1$ is an integer where the least exponent $a$ appearing in the $\varphi$-representation is odd. Then $a+1$ does not appear. Apply Galois conjugation, replacing $\varphi$ in the representation by $\overline{\varphi}$, obtaining that $n$ is also a sum of the same powers of $\overline{\varphi}$. Now since $a$ is odd, we have $\overline{\varphi}^{a}<0$. Thus the next larger positive term one could have in this new representation is $\overline{\varphi}^{a+3}$. It follows that $n\leq\overline{\varphi}^{a}+\overline{\varphi}^{a+3}+\overline{\varphi}^{a+5}+\cdots=\overline{\varphi}^{a}(1-\varphi^{-2})<0$, a contradiction.
(b) By induction on the (even) length of the string $y$. First, note that for the empty string $\varepsilon$, we can take $n=1=1.\varepsilon^{R}$. Next, for $y=10$, we can take $n=2=10.01$. Now assume that $|y|=c\geq 4$ is even, and that $y$ starts with a $1$. We distinguish between three cases according to the number of 0’s following the leading $1$ in $y$.
Case 1: The leading $1$ in $y$ is followed by at least three 0’s. We replace the prefix $1000$ in $y$ by $10$ to obtain a new string $y^{\prime}$ with $|y^{\prime}|=c-2$. Then by induction there is some $n^{\prime}$ whose $\varphi$-representation is $x^{\prime}.{y^{\prime}}^{R}$. Now consider $n=n^{\prime}-(\varphi^{c-2}+\varphi^{-(c-2)})+(\varphi^{c}+\varphi^{-c})=n^{\prime}-L_{c-2}+L_{c}$, which is clearly a positive integer. By construction (note that $\varphi^{c-2}+\varphi^{c}=\varphi^{c-1}$, which only impacts the positive part of the representation) we know that the $\varphi$-representation of $n$ has negative part $y$.
Case 2: The leading $1$ in $y$ is followed by exactly two 0’s. Since $|y|=c\geq 4$, we know that $y$ has the prefix $1001$. We replace it by $10$ to obtain a new string $y^{\prime}$ with $|y^{\prime}|=c-2$. Again, by induction there is some $n^{\prime}$ whose $\varphi$-representation is $x^{\prime}.{y^{\prime}}^{R}$. Note that $\varphi^{-c}+\varphi^{-(c-3)}-\varphi^{-(c-2)}=\varphi^{-(c-2)}$. Therefore, we can consider $n=n^{\prime}+\varphi^{-(c-2)}+\varphi^{c-2}=n^{\prime}+L_{c-2}$. This is again an integer, and by construction the $\varphi$-representation of $n$ has negative part $y$.
Case 3: The leading $1$ in $y$ is followed by exactly one 0, i.e., $y$ has the prefix $101$. Then we can replace $101$ by $1$ to obtain $y^{\prime}$, and finish as in the previous cases, simply setting $n=n^{\prime}+L_{c}$.
(c) Let $n\geq 2$ be an integer. By the first part we know that the smallest exponent in its $\varphi$-representation is even, say 2i-2i. We need to show that $L_{2i-1}<n\leq L_{2i+1}$. By Galois conjugation we have $n=\varphi^{2i}\pm\dots$, where the remaining terms have exponents $\leq 2i-2$ and all even powers occur with positive sign and all odd powers occur with negative sign. Thus we get on the one hand
@@
  • $n$ has one non-negative even exponent and no negative even exponents (which only occurs for $n=1$). If we say that the last position of $x$ is “position 11”, then this corresponds to $x$ having exactly one $1$ at an odd position (and even positions don’t matter) and $y$ having no $1$’s at even positions (and odd positions don’t matter).
@@
def thm9 "?msd_fib Ex,y \$saka(n,x,y) & \$one_even_exponent(x,y)":
@@
$\varphi^{2k}-\varphi^{-(2i+1)}=\varphi^{2k-1}+\varphi^{2k-3}+\dots+\varphi^{-(2i-1)}+\varphi^{-(2i+2)}.$
@@
Now assume conversely that the integer $x>1$ has exactly one even exponent in its $\varphi$-representation. By Theorem 8 it must be the smallest exponent, and so in particular all positive exponents are odd, and $\varphi^{0}$ does not appear. Let ${\cal O}_{\text{pos}}$ be the set of (odd) positive exponents in the $\varphi$-representation of $x$, and ${\cal A}_{\text{neg}}$ be the set of negative exponents. Consider the number
@@
eval thm10b "?msd_fib An \$sum_even_lucas(n-2) <=> \$thm10(n)":
@@
eval thm10c "?msd_fib An \$thm10(n) => Ex,y \$saka(n,x,y) & \$fibmatch(x,2)":
@@
def thm11_match "?msd_fib \$fib3(a,d) & \$fibmatch(x,d) & \$fibmatch(y,a)":
[…]
eval thm11b "?msd_fib An \$thm11(n) =>  Ex,y \$saka(n,x,y) & ((\$fibmatch(x,5)
   & \$fibmatch(x,2)) | Ea,d \$thm11_match(a,d,x,y))":
@@
Acknowledgments
We thank Michel Dekking for pointing out some flaws in a previous version.